# MiniScope: Automated UI Exploration and Privacy Inconsistency Detection of MiniApps via Two-phase Iterative Hybrid Analysis

### v1

@@ 4.1.3. Merging into MinApp Dependency Graph
To precisely model MiniApps, MiniScope combines UTG and CCFG with UDFG proposed in TaintMini (Wang et al., 2023b) to create a unified topological structure of MiniApps, called MiniApp Dependency Graph (MDG). For example, MiniScope can analyze sensitive APIs distribution and their triggering paths by querying the graph. In particular, we first formally describe UDFG following its original definition, then we define MDG accordingly.
Definition 4 (UDFG): Universal Data Flow Graph (UDFG) can be used to capture the cross-language and cross-page data flow within a MiniApp (Wang et al., 2023b). Formally, UDFG can be represented as a directed graph $G_{D}=(V_{D},E_{D},\lambda_{D})$, where $V_{D}$ is the set of data objects, $E_{D}$ is the set of edges representing sensitive data flow dependencies and propagation directions. Correspondingly, $\lambda_{D}$ is the function for tracing data flow propagation.
@@
MDG of the code snippet in the motivating example. As shown in Figure 5, to analyze the access to camera resources, MiniScope first locates the page “pages/takePhoto” (❶ Green Line) which is navigated through the <navigator> widget in “pages/index”. Then MiniScope determines the entry point of the CCFG is “takePhoto”, which is triggered by the Tap event bound to the <button> in the WXML (❷ Purple Line). Through the combination analysis of control flow and data flow, MiniScope can determine that the takePhoto function further calls wx.createCameraContext to create the context, which is then assigned to the variable “ctx”. Later, the camera is triggered to take a photo through the “ctx.takePhoto” method (❸ Pink Line).
Figure 5. MiniApp Dependency Graph (MDG) representation of the code snippet presented in Figure 2. The Black lines represent AST edges; the Green Lines represent UTG edges; the Purple Lines represent CCFG edges; and the Pink Lines represent UDFG edges.
@@ 4.2.1. Fuzzing Matching between WXML Component and UI Widget Tree
As described in Challenge#2, due to the platform-specific widgets defined in WXML (e.g., <navigator>) and the support for programmable attributes (e.g., loop rendering with wx:for), WXML components lack the unique resourceID and deterministic XPATH in the rendered UI screen, making it difficult to locate them uniquely. Therefore, to leverage the static MDG for guiding dynamic UI exploration, we propose a fuzzy matching approach to establish a robust mapping between WXML component and UI widget tree. The fuzzy matching comprises two primary strategies, as outlined below, to accurately locate widgets on the rendered UI screen.
@@
Through this fuzzy matching approach, MiniScope effectively bridges the gap between static MDG and dynamic UI exploration. By leveraging key attributes and the Levenshtein distance, MiniScope is able to accurately map WXML components to their corresponding widgets in the UI Widget Tree, overcoming the inherent challenges in locating platform-specific widgets within WXML.
@@ 4.2.2. Task-directed UI Exploration Strategy
Privacy-practice-directed DFS Exploration. Once BFS is completed, MiniScope unpacks and merge the dynamically loaded sub-package into the main package. A subsequent round of static analysis updates the MDG. In Phase Two, guided by the complete MDG, MiniScope delves into a DFS exploration focusing on privacy-related practices within the MiniApp. Similar to BFS, this exploration process involves computing GUI events on specific pages to trigger relevant callbacks. However, in contrast to BFS, the DFS exploration prioritizes completing each potential privacy practice path in a first-in, first-out order from the queue, ensuring a thorough examination of privacy practices.
Through the aforementioned two-phase task-directed UI exploration strategy, MiniScope achieves a comprehensive and systematic analysis of MiniApps, ensuring a robust and detailed understanding of MiniApp privacy practices, addressing the challenges posed by their unique features.
@@ 4.3. Privacy Inconsistency Detector
MiniScope could enable many security and software engineering tasks seamlessly. In this paper, our primary focus centers on the application of MiniScope to detect privacy inconsistency within MiniApps. Thus we employ an innovative approach to monitor the privacy practices within MiniApps by instrumenting sensitive APIs and cross validate with the statement in privacy policies to detect inconsistency.
@@ 4.3.1. Privacy Practice Monitor
In this component, MiniScope leverages the Frida framework (frida, 2023) to hook sensitive APIs. This enables us to monitor the API calls made by the MiniApp during dynamic UI exploration. When these API calls are triggered, our monitor captures crucial context information and stack traces, logging any detected statements. However, MiniApps execute their logic layer primarily through APIs that are encapsulated from native library functions via the JSBridge mechanism (Lu et al., 2020). Thus the main challenge we faced lies in the mapping between each sub-app API and the corresponding system API in Java layer. By reverse engineering WeChat, we found that invoking a MiniApp API from JavaScript triggers a native bridge function offered by the host platform, which manages the execution of the actual system APIs. In the context of WeChat MiniApps, this bridge function is located within the commonjni.AppBrandJSBridgeBinding class, specifically in the invokeCallbackHandlerI(int, java.lang.String) method. Utilizing objection (sensepost, 2023), a runtime mobile exploration toolkit powered by Frida (frida, 2023), MiniScope hooks this bridge function, checking its parameters and runtime call stack. This allows us to establish the mapping from each sub-app API to the corresponding Java layer system API (detailed in our online documentation (Anonymous, 2023)). Note that this process of constructing the WeChat-to-Android API Mapping is offline and can be reused throughout our analysis once built.
@@ 4.3.2. LLM-based Privacy Policy Analysis
Definition of Privacy Statement. As shown in Table 3, we represent privacy policies as a series of triples, i.e., (DC, SSoC, DE). Data Controller is the party responsible for determining the purposes and means of personal data processing, which can be the application itself (first party) or a third party. SSoC Verbs refer to a list of verbs that describe the Storing-Sharing-or-Collection of data. Data Entity represents any privacy information that can identify or reflect the activities of natural persons.
@@
Prompt Design. Typically, previous works (Harkous et al., 2018; Andow et al., 2019; Ling et al., 2022) rely on manually annotated corpus to train named entity recognition models (such as CRF, BiLSTM, BERT, etc.) for predicting entity labels, or by constructing data and entity dependency (DED) trees to extract tuple representations of privacy policy statements. With the recent emergence of large language models (LLMs), we aim to accomplish privacy policy extraction through fine-tuning LLMs using few-shot in-context learning. The prompt consists of three parts: (1) Task Description, (2) Few-shot Demonstrations, and (3) Query. For better understanding, we provide the following prompt example.
Task Description: Here is a table where each row has 3 items. The first item is ”privacy detail”, and the second to third items are ”subject” and ”verb” for the ”privacy detail” in the same row. Here are some sample rows in the table: Few-shot Demonstrations: ”To save photos, the developer will request your permission to access your photo album”, ”developer”, ”access”, ”photo album” Query: Please read the table and follow the above sample rows to complete the table by filling ”subject”, ”verb”, ”privacy detail” of each row: {Input Sentence}
@@ 5. Evaluation
We have implemented a prototype of MiniScope using 9,466 lines of code (LoC) with Python, excluding any third-party libraries or open-source tools. We then conducte an evaluation on MiniScope, aiming to answer the following research questions (RQs):
@@ 5.1. Experimental Setup
Dataset. To collect MiniApps, we utilize the open-source tool MiniCrawler (Zhang et al., 2021) to download MiniApp packages from the WeChat App Market. We have collected a total of 127,460 MiniApps, with 289 GB of total size. To collect privacy policies from our list of MiniApps, we designe and deploy a privacy policy crawler. It requests privacy policies from the WeChat server based on the AppID, completing the collection process within 8 hours. Due to the prevalent absence of privacy policies, we find that only 10,786 (8.4%) MiniApps have valid privacy policies.
Ground Truth. To ensure a reliable evaluation of RQ1 and RQ2, we form a benchmark by randomly sampling 150 MiniApps from the dataset. They are dynamically executed for sub-package loading and manually inspected by three experienced researchers separately, which forms the ground truth of their privacy-related practices. To ensure a thorough dynamic interaction, each researcher interacts with the MiniApp for 5 minutes or until all reachable pages/functions are fully triggered, whichever happens first.
Running Environment. Our experiments are conducted on a server running Ubuntu Linux of 22.04 version with two 64-core AMD EPYC 7713 and 256 GB RAM. The dynamic testing is performed on an Android Virtual Device (AVD) with a system version of Android 8.1.0 and API version Level 27. The version of WeChat used is 8.0.37, and the WebView kernel version is 107.0.5304.141.
@@ 5.2. Effectiveness of MiniScope (RQ1)
Effectiveness of Hybrid Analysis. We compare the performance of MiniScope with that of TaintMini (Wang et al., 2023b) on our ground truth dataset. For a comprehensive assessment, we use the following metrics: True Positives (TPs), False Positives (FPs), False Negatives (FNs), precision, recall, and F1 score.
The results of the evaluation are listed in Table 4. MiniScope significantly surpasses TaintMini in detecting 19 out of 20 sensitive APIs, offering the highest precision, recall and F1 score. The lower precision of TaintMini stems from false positives found in unreachable code, while its decreased recall is due to its omission of sub-packages and its use of a single data flow analysis. This results in TaintMini overlooking many device-specific API calls like wx.openBluetoothAdapter, which may be called only once, not involving data flow propagation. MiniScope, by incorporating page state transition analysis, callback control flow analysis, and universal data flow analysis, provides the best overall performance.
Table 4. Effectiveness of hybrid analysis using MiniScope. Prerepresents Precision%; Recrepresents Recall%; F1% represents F1-score%.
Category APIs GroundTruth TaintMini MiniScope
@@
Effectiveness of LLM-based Privacy Policy Analysis. We compare the LLM-based approach with five baseline NER models. Particularly, we utilize GPT-3.5-Turbo to perform inquiries. During this process, we keep the default configuration of GPT-3.5-Turbo, with $temperature=1$ and $top\_n=1$. To complete the experiment, we have utilized an estimation of 62 thousand tokens in total to analyze 100 privacy policies. For the other five baseline models, we train them using annotations from the CA4P-483 (Zhao et al., 2022) corpus. To ensure unbiased evaluation, we apply 10-fold cross-validation to the models, with 8 folds used for training, 1 fold for parameter tuning and optimization, and 1 fold for testing. Table 5 presents the detailed performance of these models. We observe that GPT-3.5-Turbo outperforms other baseline models significantly, indicating its high robustness and effectiveness.
@@
For hybrid analysis, MiniScope outperforms RegMatch and TaintMini on our ground truth dataset with 99.2% precision, 95.6% recall, and 97.4% F1 score. Similarly, in privacy policy analysis, our LLM-based method surpasses existing NER models with 98.5% precision, 94.5% recall, and 96.5% F1 score.

### v3

@@ 4.1.3. Merging into MinApp Dependency Graph
To precisely model MiniApps, MiniScope combines UTG and CCFG with UDFG proposed in TaintMini (Wang et al., 2023a) to create a unified topological structure of MiniApps, called MiniApp Dependency Graph (MDG). For example, MiniScope can analyze sensitive APIs distribution and their triggering paths by querying the graph. In particular, we first formally describe UDFG following its original definition, then we define MDG accordingly.
Definition 4 (UDFG): Universal Data Flow Graph (UDFG) can be used to capture the cross-language and cross-page data flow within a MiniApp (Wang et al., 2023a). Formally, UDFG can be represented as a directed graph $G_{D}=(V_{D},E_{D},\lambda_{D})$, where $V_{D}$ is the set of data objects, $E_{D}$ is the set of edges representing sensitive data flow dependencies and propagation directions. Correspondingly, $\lambda_{D}$ is the function for tracing data flow propagation.
@@
MDG of the code snippet in the motivating example. As shown in Figure 5, the MDG captures multiple layers of interactions to analyze key behaviors, such as accessing camera resources. First, MiniScope identifies the navigation from ‘‘pages/myInfo’’ to ‘‘pages/takePhoto’’ via the <navigator> component, where the path is dynamically bound to the JavaScript variable takePhotoPath (❷ Green Line). This demonstrates dynamic data binding between the logical and rendering layers. Next, MiniScope identifies the event handler onShutterTap (❸ Purple Line), which is bound to the Tap event of the <button> widget in ‘‘pages/takePhoto’’. Unlike statically defined callbacks, onShutterTap is dynamically loaded by the init function in pages/util/util.js, which is invoked during the onReady lifecycle method. Through context binding (❹ Pink Line), MiniScope can identify event handler callbacks defined across pages. Through combined control and data flow analysis, MiniScope determines that the onShutterTap function invokes wx.createCameraContext to create a camera context, which is assigned to the variable ctx. Subsequently, the camera is triggered to take a photo using the ctx.takePhoto method. The captured image file path is then transmitted back to the page via pageContext.setData, dynamically updating the imagePath property.
Figure 5. MDG representation of the code snippet presented in Figure 2. The Black lines represent AST edges; the Green Lines represent UTG edges; the Purple Lines represent CCFG edges; and the Pink Lines represent UDFG edges. For simplicity, some AST nodes and edges have been omitted.
@@ 4.2.1. Fuzzing Matching between WXML Component and UI Widget Tree
As described in Challenge#2, due to the platform-specific widgets defined in WXML (e.g., <navigator>) and the support for programmable attributes (e.g., loop rendering with wx:for), WXML components lack the unique resourceID and deterministic XPATH in the rendered UI screen, making it difficult to locate them uniquely. Therefore, to leverage the static MDG for guiding dynamic UI exploration, we propose a fuzzy matching approach to establish a robust mapping between the WXML component and the UI widget tree. The fuzzy matching comprises two primary strategies, as outlined below, to accurately locate widgets on the rendered UI screen.
@@
Through this fuzzy matching approach, MiniScope effectively bridges the gap between static MDG and dynamic UI exploration. By leveraging key attributes and the Levenshtein distance, MiniScope is able to accurately map WXML components to their corresponding widgets in the UI Widget Tree, overcoming the inherent challenges in locating platform-specific widgets.
@@ 4.2.2. Task-directed UI Exploration Strategy
Privacy-practice-directed DFS Exploration. Once BFS is completed, MiniScope unpacks and merge the dynamically loaded sub-package into the main package. A subsequent round of static analysis updates the MDG. In Phase Two, guided by the complete MDG, MiniScope delves into a DFS exploration focusing on privacy-related practices within the MiniApp. Similar to BFS, this exploration process involves computing GUI events on specific pages to trigger relevant callbacks. However, in contrast to BFS, the DFS exploration prioritizes completing each potential privacy practice in a first-in, first-out order from the queue, ensuring a thorough examination.
Through the aforementioned two-phase task-directed UI exploration strategy, MiniScope achieves a comprehensive and systematic analysis of MiniApps, ensuring a robust and detailed understanding of MiniApp privacy practices, and addressing the challenges posed by their unique features.
@@ 4.3. Privacy Inconsistency Detector
MiniScope could enable many security and software engineering tasks seamlessly. In this paper, our primary focus centers on the application of MiniScope to detect privacy inconsistency within MiniApps. Thus we monitor the privacy practices within MiniApps by instrumenting sensitive APIs and cross-validate with the statement in privacy policies to detect inconsistency.
@@ 4.3.1. Privacy Practice Monitor
In this component, MiniScope leverages the Frida framework (frida, 2023) to hook sensitive APIs. This enables us to monitor the API calls made by the MiniApp during dynamic UI exploration. When these API calls are triggered, our monitor captures crucial context information and stack traces, logging any detected statements. A key challenge arises from the architecture of MiniApps, where their logic layer primarily interacts with underlying functionalities through APIs encapsulated via the JSBridge mechanism (Lu et al., 2020). This encapsulation abstracts the connection between MiniApp APIs and the native system APIs, making it non-trivial to establish a direct mapping. To address this, we drew insights from the prior work (Lu et al., 2020) and replicated their approach as part of our analysis. Specifically, we started by reverse engineering the WeChat platform to analyze its JSBridge mechanism. We found that invoking a MiniApp API from JavaScript triggers a native bridge function offered by the host platform, which manages the execution of the actual system APIs. In the context of WeChat MiniApps, this bridge function is located within the commonjni.AppBrandJSBridgeBinding class, specifically in the invokeCallbackHandlerI(int, java.lang.String) method. Utilizing objection (sensepost, 2023), a runtime mobile exploration toolkit powered by Frida (frida, 2023), MiniScope hooks this bridge function, checking its parameters and runtime call stack. This allows us to establish the mapping from each sub-app API to the corresponding Java layer system API (detailed in our online documentation (Anonymous, 2023)). Note that this process of constructing the WeChat-to-Android API Mapping is offline and can be reused throughout our analysis once built.
@@ 4.3.2. LLM-based Privacy Policy Analysis
Definition of Privacy Statement. As shown in Table 3, we represent privacy policies as a series of triples, i.e., (DC, SSoC, DE). Data Controller is the party responsible for determining the purposes and means of personal data processing, which can be the application itself (first party) or a third party. SSoC Verbs refer to a list of verbs that describe the Storing-Sharing-or-Collection of data. Data Entity represents any privacy information or sensitive permission.
@@
Few-shot Prompt Design. Typically, previous works (Harkous et al., 2018; Andow et al., 2019; Ling et al., 2022) rely on manually annotated corpus to train named entity recognition models (such as CRF, BiLSTM, BERT, etc.) for predicting entity labels, or by constructing data and entity dependency (DED) trees to extract tuple representations of privacy policy statements. With the recent emergence of large language models (LLMs), we aim to leverage the power of LLMs to extract fine-grained privacy statements from the policy. Instead of training the model from scratch, we utilize few-shot learning by providing the LLM with a small set of annotated examples (few-shot demonstrations). These examples were carefully selected to cover a range of typical privacy policy statements involving different data controllers, actions, and data entities. Specifically, the prompt constructed consists of three parts: (1) Task Description, (2) Few-shot Demonstrations, and (3) Query. For better understanding, we provide the following prompt example.
The Prompt Template Task Description: Here is a table where each row has 4 items. The first item is “privacy statement”, and the second to fourth items are “subject”, “verb”, and “object” extracted from the “privacy statement”. Here are some samples: Few-shot Demonstrations: Demo#1: “To save photos, the developer will request your permission to access your photo album”, “developer”, “access”, “photo album”. Demo#2: “In order to help you become our member, the developer will collect your WeChat nickname and avatar after obtaining your express consent”, “developer”, “collect”, “WeChat nickname and avatar”. …… Query: Please read the table and follow the above sample rows to complete the table by filling in “subject”, “verb”, “object” of each row: {Input Privacy Statement}.
@@ 5. Evaluation
We have implemented a prototype of MiniScope using 9,466 lines of code (LoC) with Python, excluding any third-party libraries or open-source tools. We then conduct an evaluation on MiniScope, aiming to answer the following research questions (RQs):
@@ 5.1. Experimental Setup
Dataset. To collect MiniApps, we utilize the open-source tool MiniCrawler (Zhang et al., 2021) to download MiniApp packages from the WeChat App Market. We have collected a total of 127,460 MiniApps, with 289 GB of total size. To collect privacy policies from our list of MiniApps, we design and deploy a privacy policy crawler. It requests privacy policies from the WeChat server based on the AppID, completing the collection process within 8 hours. Due to the prevalent absence of privacy policies, we find that only 10,786 (8.4%) MiniApps have valid privacy policies.
Ground Truth. To ensure a reliable evaluation of RQ1 and RQ2, we carefully curate a comprehensive ground truth dataset by randomly sampling 100 MiniApps, which consists of two parts. To evaluate the performance of hybrid analysis, three experienced researchers separately interacted with these MiniApps for sub-package loading and manually inspected the privacy collected by the MiniApp, which forms the ground truth of their privacy-related practices. To ensure a thorough dynamic interaction, each researcher interacts with the MiniApp for 5 minutes or until all reachable pages/functions are fully triggered, whichever happens first. To evaluate the performance of LLM-based privacy policy analysis, we create a benchmark dataset by manually annotating 100 privacy policies associated with the sampled MiniApps. Specifically, three researchers independently read through each privacy policy and manually labeled the privacy practices mentioned in the text. This carefully curated benchmark, comprising 674 manually annotated privacy statements, serves as the ground truth for evaluating the performance of our privacy policy analysis techniques.
Running Environment. Our experiments are conducted on a server running Ubuntu Linux of 22.04 version with two 64-core AMD EPYC 7713 and 256 GB RAM. The static analysis leverages the server’s computational capacity by utilizing 128 threads, enabling high parallelism for efficient processing of the large MiniApp dataset. The dynamic testing is performed on 16 Android Virtual Devices (AVDs) running in parallel, each configured with a system version of Android 8.1.0 and API Level 27. The version of WeChat used is 8.0.37, and the WebView kernel version is 107.0.5304.141.
@@ 5.2. Effectiveness of MiniScope (RQ1)
Table 4. Effectiveness of hybrid analysis using MiniScope. GT represents the Ground Truth; Prerepresents Precision%; Recrepresents Recall%; F1% represents F1-score%.
Category APIs GT TaintMini MiniScope
@@
Effectiveness of Hybrid Analysis. We compare the performance of MiniScope with that of TaintMini (Wang et al., 2023a) on our ground truth dataset. We use the following metrics: True Positives (TPs), False Positives (FPs), False Negatives (FNs), precision, recall, and F1 score. The results of the evaluation are listed in Table 4. MiniScope significantly surpasses TaintMini in detecting 19 out of 20 sensitive APIs, offering the highest F1 score. The lower precision of TaintMini stems from FPs found in unreachable code (unused function and orphaned pages), while its decreased recall is due to its omission of sub-packages and its use of a single data flow analysis. For example, TaintMini only includes callbacks in its taint analysis when they are involved in data flows from WXML to JavaScript. This leads to TaintMini failing to determine the entry points of callback control flow in many cases, resulting in false positives related to unused functions. Furthermore, TaintMini overlooks many device-specific API calls like wx.openBluetoothAdapter, which may be called only once, not involving data flow propagation. MiniScope, by incorporating UTG, CCFG, and UDFG, provides the best overall performance. However, it is important to note that MiniScope offers the lower precision compared to TaintMini for 4 specific APIs, which can be attributed to the path insensitivity of MiniScope’s broader analysis scope. While subpackage loading provides a broader analysis scope, allowing MiniScope to detect more TPs, its path insensitivity leads to the merging of conditional branches, thereby introducing more FPs and affecting the precision metric. Although MiniScope effectively reduces most FPs caused by unused functions and orphaned pages, it cannot completely eliminate all of them due to its inherent path insensitivity.
Time Efficiency of MiniScope. In addition, we analyze the time efficiency of MiniScope to further evaluate its practicality. The analysis time is broken down into four components. First, the static analysis of the main package requires an average of 29.63 seconds per MiniApp, which involves analyzing the JavaScript and WXML code and extracting page transition, callback control flow, and data flow information. Second, sub-package-directed BFS UI exploration, which dynamically explores MiniApp pages based on the structure of sub-packages, takes approximately 64.97 seconds on average, depending on the number and depth of the sub-packages. Third, the complete package static analysis, which integrates the main package and sub-packages, requires around 36.54 seconds per MiniApp. Finally, privacy-practice-directed DFS UI exploration, which focuses on exploring privacy-relevant behavior and detecting sensitive API usage, is the most time-intensive phase, averaging 222.02 seconds per MiniApp. It is worth noting that MiniScope does not guarantee full UI coverage during testing, as covering all pages is unnecessary to achieve the testing objectives. These results indicate that the UI exploration phases, including sub-package-directed and privacy-practice-directed exploration, are the most time-consuming due to the complexity of dynamic exploration. Despite this, the time costs remain reasonable given the comprehensive analysis.
Effectiveness of LLM-based Privacy Policy Analysis. We compare the performance of our LLM-based approach with five baseline NER models on the ground truth dataset of 100 manually annotated privacy policies. Specifically, we leverage GPT-3.5-Turbo to perform inquiries and extract privacy statements. During this process, we keep the default configuration of GPT-3.5-Turbo, with $temperature=1$ and $top\_n=1$. To complete the experiment, we have utilized an estimation of 62k tokens in total to analyze 100 privacy policies. For the other five baseline models, we train them using annotations from the CA4P-483 (Zhao et al., 2022) corpus. To ensure unbiased evaluation, we apply 10-fold cross-validation to the models, with 8 folds used for training, 1 fold for parameter tuning and optimization, and 1 fold for testing. We then leverage these baseline models to generate privacy statement annotations for the ground truth dataset. The performance of the LLM-based approach and the five baseline models is evaluated using standard metrics: Precision, Recall, and F1-score. The detailed performance results are presented in Table 5. We observe that GPT-3.5-Turbo outperforms all five baseline models significantly, indicating its high robustness and effectiveness in extracting privacy practices from textual privacy policies.
@@
ANSWER to RQ1 For hybrid analysis, MiniScope outperforms TaintMini on our ground truth dataset with 99.2% precision, 95.6% recall, and 97.4% F1 score. Similarly, in privacy policy analysis, our LLM-based method surpasses existing NER models with 98.5% precision, 94.5% recall, and 96.5% F1 score.